The total vapor pressure above a binary solution of components A and B is found to be 156.7 torr. Assuming ideal behavior, what is the mole fraction of A if PA* = 233.3 torr , PB* = 66.1 torr
Answer #1

Let mole fraction of A be x, then mole fraction of B = 1-x

Use:
P = PA*xA + PB*xB
156.7 = 233.3*x + 66.1*(1-x)
156.7 = 233.3*x + 66.1 - 66.1*x
156.7 = 167.2*x + 66.1
167.2*x = 90.6
x = 0.542

Answer: 0.542